OpenStudy (anonymous):

In math class, the girl to boy ratio is 3 to 9. If there are 15 girls, how many boys are there?

OpenStudy (adamaero):

http://www.purplemath.com/modules/ratio.htm

OpenStudy (studygurl14):

Set up a proportion of girls to boys \(\large\frac{3}{9}=\frac{15}{b}\) Solve for b

OpenStudy (anonymous):

15/3=5 so 9*5=45boys

OpenStudy (anonymous):

thx